The current share price of a stock is $95.00. The stock will pay an annual dividend of $7.00 in the following 3 years. The first dividend will be paid one year from now. The continuous compounding interest rate is 4%. Determine the forward price of the forward contract that matures in 3 years. Round your answer to the nearest dollar.
Answer by ikleyn(52785)

.
The current share price of a stock is $95.00.
The stock will pay an annual dividend of $7.00 in the following 3 years.
The first dividend will be paid one year from now.
The continuous compounding interest rate is 4%.
Determine the forward price of the forward contract that matures in 3 years.
Round your answer to the nearest dollar.
~~~~~~~~~~~~~~~~

Make a Table


    Year    make                   pay            Balance
            compounding            dividend


      1     95*2.71828^0.04        98.88 - 7      = 91.88
            = 98.88


      2     91.88*2.71828^0.04     95.63 - 7      = 88.63
            = 95.63


      3     88.63*2.71828^0.04     92.25 - 7      = 85.25
            = 92.25


ANSWER.  After 3 years, the stock is $85.25; the total dividend paid in 3 years is $21.

         The total price is  $85.25 + $21 = $106.21.


         Round as you want.